Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 337     Entry time: 09 Sep 2023, 07:00
Author: Erik Brambrink 
Group: HED 
Subject: Shift summary 
  • Data taken from 11 to cryo failure
  • Beam conditions stable
  • Timing scan for different energy
  • Operation stopped after cryo failure
  • Cryo jet was heated up then
  • Improvements of optical diagnostics (far field monitor)
ELOG V3.1.4-7c3fd00